What is the distance between the points (3,2) and (7,10)
Answer by checkley71(8403)

THESE POINTS FORM A RIGHT TRIANGLE WITH SIDES=10-2=8 FOR ONE SIDE & 7-3=4 FOR THE OTHER SIDE. THUS USING THE PYTHAGOREAN THEORUM WE GET
4^2+8^2=D^2
16+64=D^2
D^2=80
D=SQRT80
D=8.94 ANSWER FOR THE DISTANCE BETWEEN THESE POINTS.